OpenStudy (anonymous):

Suppose that a small pump can empty a swimming pool in 70 hours and that a large pump can empty a pool in 30. Working together, how long will it take to empty the pool?

OpenStudy (anonymous):

Is it 21?

OpenStudy (anonymous):

Set it up the same way as the other one. S = small pump flowrate = 1/70 L = large pump flowrate = 1/30 (1/70 + 1/30)*t = 1 solve for t

OpenStudy (anonymous):

Yes! =D 21